OpenStudy (jpsmarinho):

Lim (1- cosx)/x when x goes to 0 How to solve this without L'Hospital?

OpenStudy (experimentx):

\[ 1 - \cos x = 2 \sin^2 (x/2) \] use the property that sin x/x = 1

OpenStudy (jpsmarinho):

I can substitute 1- cosx per something? I thinking in something related to the trigonometric identity to obtain sinx

OpenStudy (jpsmarinho):

Ah I multiply the num and denominator for the conjugate of the the numerator ^^. Thanks!!

OpenStudy (experimentx):

 0x1 = 0